Just before striking the ground, a 2.0-kg mass has 400 J of KE. If friction can be ignored, from what height was it dropped ?

1
Expert's answer
2022-03-28T11:02:20-0400

The law of conservation of energy says

"PE_i=KE_f"

"mgh_i=KE_f"

Hence

"h_i=\frac{KE_f}{mg}=\frac{400}{2.0*9.8}=20\:\rm m"